OpenStudy (anonymous):

Find the polynomial function of lowest degree with rational coefficients thaat has the given numbers as some of its zeros. 4+i,2

OpenStudy (anonymous):

other zero is \[4-i\] so your job is to multiply \[(x-2)(x-(4+i))(x-(4-i))\]

OpenStudy (anonymous):

which is not as hard as it might look. do you know how to do it? the snap way i mean?

OpenStudy (anonymous):

No i get stuck after the first part of solving it

OpenStudy (anonymous):

i can get x\[x ^{3^{?}}\]

OpenStudy (anonymous):

ok first off we have to multiply \[(x-(4+i))(x-(4-i))\] and it is pretty straight forward. the "last" term, that is the constant, will be \[(4+i)(4-i)\] clear?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and it is always true that \[(a+bi)(a-bi)=a^2+b^2\] so you get a constant of \[4^2+1^2=17\] notice we do not compute with any i's

OpenStudy (anonymous):

the "outer , inner" will be \[-(4+i)x -(4-i)x=8x\]

OpenStudy (anonymous):

so we have \[(x-(4+i))(x-(4-i))=x^2-8x+17\] done

OpenStudy (anonymous):

now you have to multiply this by x -2 but that is straghtforward

OpenStudy (anonymous):

you should convince yourself that \[(x-(a+bi))(x-(a-bi))=x^2-2ax+a^2+b^2\] so it is not hard

OpenStudy (anonymous):

your "final answer" will be \[(x-2)(x^2-8x+17)\] whatever that is

OpenStudy (anonymous):

after solving that equation I will have my answer?

OpenStudy (anonymous):

yes. here it is, and you can even see if you scroll down that we have the correct zeros http://www.wolframalpha.com/input/?i=%28x-2%29%28x^2-8x%2B17%29

OpenStudy (anonymous):

real root is 2, complex root is 4 + i and its conjugate 4 - i

OpenStudy (anonymous):

okay I beileve I got it

OpenStudy (anonymous):

Thank you.